OpenStudy (anonymous):

9x^2-38x+8 a. fins two numbers whose product is 9*8=72 and whose sum is -38. b. write -38x using the factors from part (a).

OpenStudy (mathmate):

Make a table of the candidates and check if they fit the criteria A B A+B AB -1 -72 -73 72 -2 -36 -38 72 -3 -24 -27 72 -4 -18 -22 72 -6 -12 -18 72 .... We notice that -2 and -36 fit the given requirements of sum=-38, product=72. So 9x^2-38x+8 = 9x^2- 36x -2x +8 = 9x(x-4) - 2(x-4) = (9x-2)(x-4) Notice that 9*4=36, 2*1=2, and -36-2=-38

OpenStudy (mathmate):

Note: the idea is that as you do more of these exercises, the tabular part can be done mentally (that's why knowing the times table is important in math).